OpenStudy (klepez):

What is the excluded value of the rational expression ? (1 point) –1 4 –7 None of the above

OpenStudy (klepez):

\[\frac{ x -4}{6 }\div \frac{ x+1 }{ 7 }\] is the equattion

OpenStudy (nubeer):

u have to put denominator =0 in this case ur denominator is (x+1)/7 =0 solve it

OpenStudy (abb0t):

flip the right fraction so that the 7 is in the numerator and x+1 is in denominator and multiply/.

OpenStudy (precal):

just like abb0t stated above, now look at the denominator and see what number x can not be